OpenStudy (anonymous):

Solve for V r = sqrt(3V/πh) I squared both sides to remove the radical, but honestly that is all my brain will process. I need help, again. :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[r^2=\frac{3V}{\pi h}\]

OpenStudy (anonymous):

you are looking for \(V\) right?

OpenStudy (anonymous):

I mean solve for V. there should be a space after before the r yes

OpenStudy (anonymous):

once more suppose you wanted to solve \[18=\frac{3V}{35}\] what would you do?

OpenStudy (anonymous):

uhhhmm no clue

OpenStudy (anonymous):

i don't mean to torture you with this you would 1) multiply by 35, then 2) divide by 3

OpenStudy (anonymous):

do the same think for \[r^2=\frac{3V}{\pi h}\] multiply by \(\pi h\) to get \[\pi hr^2=3V\] then divide by 3 to get \[\frac{\pi hr^2}{3}=V\]

OpenStudy (anonymous):

Boom! I really appreciate your help. I got myself in over my head with this class.